OpenStudy (anonymous):

help please.... x^3-3x^2-4x+12=0

OpenStudy (georgesyip):

x(x^2-3x-4)=-12

OpenStudy (georgesyip):

not sure i'm right or not... I will try again...

OpenStudy (georgesyip):

x^3-3x^2-4x+12=0 x^3-3x^2-4x=-12 x(X^2-3x-4)=-12 x(x-3)(x+1)=-12

OpenStudy (georgesyip):

this is what i get hope it helps...

OpenStudy (anonymous):

hmm not sure...the answers are 2,-2,3 and i am not getting that not sure what i am doing wrong?